can someone please please help me simplify?
1+3/7
over
3-1/3

Answer by stanbon(75887)

1+3/7
over
3-1/3
----------
[1+(3/7)]/[3-(1/3)]
=[(7/7)+(3/7)]/[(9/3)-(1/3)]
=[(10/3)]/[(8/3)]
=(10/3)*(3/8)
Cancel the 3 factors to get:
=10/8
Reduce to get:
=5/4
=============
Cheers,
Stan H.